If a car ca accelerate at 4.3m/s[sup]2[/sup], how far will it have to travel from rest to reach a speed of 28 m/s?
1
Expert's answer
2011-10-13T08:24:43-0400
a=4.3 m/s2
V=28 m/s
S-?

S=S0+V0t+at2/2
V=V0+at
S0=0, V0=0 => S=at2/2
V=at => t=V/a
S=a/2*V2/a2=V2/2a
S=282/2/4.3=91,16m